OpenStudy (anonymous):

find the derivative of : \[y=(5x+\sin^{-1}3x)^{{e^x}}\]

OpenStudy (amistre64):

implicit or partial?

OpenStudy (amistre64):

implicit with respect to t is: \[y=(5+\sin^{-1}(3x))^{{e^X}}\] \[y'=x' e^x(5+\sin^{-1}(3x))^{{e^X-1}} \frac{3x'} {\sqrt{1-(3x)^2}}\] maybe? that e^x has me concerned

OpenStudy (anonymous):

i know me too but it looks right

OpenStudy (amistre64):

wolfram says its good, except for that "3" i put up there

OpenStudy (amistre64):

if its implicit with respect to x; then x'=1

OpenStudy (anonymous):

ahhh yes wolframalpha i should have plugged and chugged

OpenStudy (anonymous):

just the variable in the exponent was bothering me, but i understand now, thanks for the help

OpenStudy (amistre64):

youre welcome :)